the wheel on a truck are turning at 430 revolution per minutes as the truck travels 40 miles per hour. What is the radius of the wheel?
Answer by lwsshak3(11628)

the wheel on a truck are turning at 430 revolution per minutes as the truck travels 40 miles per hour. What is the radius of the wheel?
**
The problem is to find the distance the wheels travel per revolution.
This distance=circumference of the wheel from which you can calculate the radius.
..
2πr=(40 mi/hr)/(400 rev/min)
=40 mi/hr*min/400 rev*hr/60 min*5280 ft/mi
mi, hr, min & rev, cancel out, leaving:
40*5280 ft/400*60=8.8 ft
2πr=8.8ft
r=8.8/2π=1.4 ft (radius of wheel)